OpenStudy (kunaliiserb):

In winter, a lake of water is covered initially by a uniform layer of ice of thickness 1 cm. The air temperature at the surface of the ice is −0.5degree C. Estimate the rate at which the layer of ice begins to thicken, assuming that the temperature of the water just below the ice is 0 degree C. You can also assume steady-state conditions and ignore convection. The temperature of the water at the bottom of the lake, depth 1 m, is maintained at 2degree C. Find the thickness of ice that will eventually be formed. The thermal conductivity of ice is 2.3Wm^-1K^-1 and of water is 0.56Wm^-1K^-1.

OpenStudy (osprey):

Brrrrrrr ! @IrishBoy Just scanned this one .. looks pretty juicy

OpenStudy (irishboy123):

if i were to have a crack at this, i'd start here and look at the start position and also the ultimate end steady state when \(\dfrac{Q}{t} = \dot Q\) is the same from the bottom of the pond all the way to the outside air......for a volume element of unit cross section area, as there's a ton of symmetry on here.  So just like modelling heat flow on 2 solid, connected rods. In terms of the time calculation, i think we should look at the latent heat for ice forming. The idea would be to say that \(\dot Q = \dot m L\) As is explained here....  such that \(\dot m = \dot x \cdot 1 = \dot x\) as we are using a unit x-sect area in our little physical model This resource looks a good refresher.... https://en.wikipedia.org/wiki/Latent_heat you could go at this with the continuity equation and Fourier's heat law but that's just nuts, IMHO. this approach compresses it into 1-D and it looks linear and do-able, ....though i haven't tried it myself :-|

OpenStudy (irishboy123):

ooops \(\dot m =\color{red}{\rho} ( \dot x \cdot 1) = \color{red}{\rho} \dot x\), maybe assuming constant density as it's an estimate... :$